OpenStudy (anonymous):

3(q-5)=2(q+5) Answer? ( its a full Equation i think )

OpenStudy (katrinakaif):

Well, first step, distribute

OpenStudy (anonymous):

Sorry i am Weak in this stuff and yes Dis. Then Combind like terms

OpenStudy (katrinakaif):

Lets distribute 3 to (q - 5)

OpenStudy (katrinakaif):

3q - 15

OpenStudy (katrinakaif):

Then distribute 2 to (q + 5) 2q + 10

OpenStudy (katrinakaif):

Now lets put it all together: 3q - 15 = 2q + 10

OpenStudy (katrinakaif):

Do the same as we did on the previous one. Constants to one side, variables on the other

OpenStudy (katrinakaif):

To do so, lets substract 2q from 3q : 3q - 2q - 15 = 10 1q - 15 = 10 Now solve as a linear equation 1q = 10 + 15 1q = 25 q = 25/1 q = 25

OpenStudy (anonymous):

Thank you :) again

OpenStudy (katrinakaif):

Anytime =)

OpenStudy (anonymous):

Lol your nice! Do you do this for a living or something ?

OpenStudy (katrinakaif):

Hahaha, no no, I indeed have a life other than being on Openstudy. It gradually becomes quite addicting and an everyday routine. Users here are not payed, simply an international community here to help ;)

OpenStudy (anonymous):

You Seem not from the United States

OpenStudy (katrinakaif):

I am not originally from the US by any means, but I do live there :P

OpenStudy (anonymous):

lol i am from the UK in usa Now lol this site is very slow!

OpenStudy (anonymous):

Havent you noticed?

OpenStudy (katrinakaif):

Using Chrome helps ;) I use Linux with a FireFox browser, its working quite quickly for me :D

OpenStudy (anonymous):

Lol i am on a Mac 2012 Lol I love money dont you ?

OpenStudy (katrinakaif):

Money doesn't buy happiness does it?

OpenStudy (katrinakaif):

*Everlasting happiness

OpenStudy (anonymous):

Well it buys me what ever i want lol